### v1

@@ 3.3 Model Scaling Strategy
The proposed CMAFNet architecture adopts a unified scaling paradigm that enables systematic adaptation across diverse deployment scenarios without modifying the underlying network topology. Following the compound scaling methodology established in modern detection frameworks, the architecture defines five capacity configurations—nano (n), small (s), medium (m), large (l), and extra-large (x)—through coordinated adjustment of two scaling coefficients: depth_multiple, which governs the repetition count of stackable blocks within each stage, and width_multiple, which controls the channel dimensionality at each layer. An additional max_channels parameter imposes an upper bound on deep-layer channel expansion to prevent computational overhead from scaling superlinearly in dual-branch fusion architectures. This parameterization produces a family of models sharing identical architectural topology and module composition, differing only in representational capacity and computational cost.
[…]
@@
The rationale for incorporating the max_channels constraint merits particular attention in the context of dual-branch RGB-D architectures. Unlike single-stream detectors where channel expansion propagates uniformly through the network, the proposed fusion strategy concatenates features from parallel branches at $P_{4}$ and $P_{5}$ scales, effectively doubling the channel count at fusion points before subsequent compression. Without explicit bounds, applying large width multipliers would cause deep-layer channels to exceed practical memory limits, particularly at the $P_{5}$ fusion stage where the Contextual Semantic Integration Framework performs global attention operations with quadratic memory complexity in channel dimension. The max_channels parameter set to 512 for medium through extra-large configurations ensures that fusion-stage channels remain tractable while permitting aggressive width expansion at shallower scales where cross-modal concatenation has not yet occurred.
@@
The selection of an appropriate model scale should be guided by a hierarchical decision framework that prioritizes task-side constraints over deployment-side constraints when both are applicable. Transmission-line defect detection presents a canonical example of task-driven scaling requirements: with 94.51% of annotated instances in the TL-RGBD dataset classified as small objects under COCO criteria, the detection task demands high-capacity models capable of preserving fine-grained features through deep pyramid pathways while accommodating the expanded representational requirements of cross-modal fusion. Small targets impose stringent demands on feature resolution and semantic granularity—characteristics that lightweight models struggle to maintain as features propagate through successive downsampling and fusion operations. Consequently, when the application domain exhibits predominant small-object distributions or requires high localization precision for subtle defects, the recommended configuration corresponds to large or extra-large scales regardless of available computational resources.
@@
When task-side constraints do not unambiguously dictate capacity requirements—for instance, in scenarios involving mixed object scales or moderate precision tolerances—the deployment platform becomes the primary determinant of model selection. Edge devices such as UAV-mounted inference units or embedded processors impose severe constraints on power consumption, memory bandwidth, and computational throughput, necessitating the nano configuration to achieve stable real-time inference despite moderate accuracy reduction. Mobile computing platforms including portable inspection terminals and field workstations offer intermediate computational capacity, rendering the small configuration appropriate for balancing responsiveness with detection fidelity. Desktop GPUs and server-class accelerators provide sufficient resources to support medium through extra-large configurations, with the specific choice determined by latency requirements and offline versus real-time operational modes. For batch processing of archived inspection imagery where throughput rather than latency governs efficiency, the extra-large configuration maximizes detection accuracy without deployment constraints.
@@ 3.7 Modality Ablation Study
Table 4 reports the modality ablation results. Using RGB images alone significantly outperforms the depth-only setting, indicating that appearance cues remain the dominant source for transmission-line defect detection. However, incorporating depth information via the proposed CMAF-Net consistently improves all accuracy metrics. In particular, mAP50 increases from 0.215 (RGB-only) to 0.243, and APs rises from 0.080 to 0.087, demonstrating that depth provides complementary geometric cues, especially for small or low-contrast defects. These results validate both the effectiveness of depth modality and the necessity of cross-modal alignment and fusion.
Table 4: Modality ablation of CMAFNet on the TL-RGBD dataset, highlighting the contribution of RGB, depth, and RGB–D fusion inputs to detection capability.

### v2

@@ 3.3 Model Scaling Strategy
CMAFNet adopts a unified scaling paradigm that enables systematic adaptation across diverse deployment scenarios without modifying the underlying network topology. Following the compound scaling methodology established in modern detection frameworks, the architecture defines five capacity configurations—nano (n), small (s), medium (m), large (l), and extra-large (x)—through coordinated adjustment of three scaling coefficients. The depth_multiple coefficient governs the repetition count of stackable blocks within each stage, the width_multiple coefficient controls channel dimensionality at each layer, and the max_channels parameter imposes an upper bound on deep-layer channel expansion to prevent computational overhead from scaling superlinearly in dual-branch fusion architectures. This parameterization produces a family of models sharing identical architectural topology and module composition, differing only in representational capacity and computational cost.
Table 2 summarizes the scaling coefficients and resulting computational characteristics for each configuration. The five configurations establish a progressive capacity gradient spanning from 4.9M parameters at the nano scale to 100.7M parameters at the extra-large scale, while maintaining inference speeds ranging from 227.6 FPS to 62.6 FPS on standard GPU hardware. The transitions between adjacent scales embody distinct capacity expansion strategies. The nano-to-small transition increases channel width while preserving depth, thereby expanding feature dimensionality for richer local representations. The small-to-medium transition further widens channels to baseline capacity while introducing deep-layer channel capping to constrain fusion-stage computation. The medium-to-large transition restores full block repetition counts to enhance hierarchical abstraction depth. The large-to-extra-large transition extends mid-to-high-level channel width beyond baseline under continued deep-layer capping, maximizing cross-modal fusion capacity.
Table 2: Scaling configurations and computational characteristics of CMAFNet variants.
Scale Depth Width Max Channels Params (M) GFLOPs
Nano (n) 0.50 0.25 1024 4.9 12.4
Small (s) 0.50 0.50 1024 14.6 38.9
Medium (m) 0.50 1.00 512 36.9 108.9
Large (l) 1.00 1.00 512 62.9 196.6
Extra-large (x) 1.00 1.50 512 100.7 337.4
@@
The rationale for incorporating the max_channels constraint merits attention in the context of dual-branch RGB-D architectures. Unlike single-stream detectors where channel expansion propagates uniformly through the network, the proposed fusion strategy concatenates features from parallel branches at $P_{4}$ and $P_{5}$ scales, effectively doubling the channel count at fusion points before subsequent compression. Without explicit bounds, applying large width multipliers would cause deep-layer channels to exceed practical memory limits, particularly at the $P_{5}$ fusion stage where the Contextual Semantic Integration Framework performs global attention operations with quadratic memory complexity in channel dimension. Setting max_channels to 512 for medium through extra-large configurations ensures that fusion-stage channels remain tractable while permitting aggressive width expansion at shallower scales where cross-modal concatenation has not yet occurred.
@@
The selection of an appropriate model scale should be guided by a hierarchical decision framework that prioritizes task-side constraints over deployment-side constraints when both are applicable. Transmission-line defect detection presents a canonical example of task-driven scaling requirements: with 94.51% of annotated instances in the TL-RGBD dataset classified as small objects under COCO criteria, the detection task demands high-capacity models capable of preserving fine-grained features through deep pyramid pathways while accommodating the expanded representational requirements of cross-modal fusion. Small targets impose stringent demands on feature resolution and semantic granularity—characteristics that lightweight models struggle to maintain as features propagate through successive downsampling and fusion operations. When the application domain exhibits predominant small-object distributions or requires high localization precision for subtle defects, large or extra-large configurations are recommended regardless of available computational resources.
@@
When task-side constraints do not unambiguously dictate capacity requirements—for instance, in scenarios involving mixed object scales or moderate precision tolerances—the deployment platform becomes the primary determinant of model selection. Edge devices such as UAV-mounted inference units or embedded processors impose severe constraints on power consumption, memory bandwidth, and computational throughput, necessitating the nano configuration to achieve stable real-time inference despite moderate accuracy reduction. Mobile computing platforms including portable inspection terminals and field workstations offer intermediate computational capacity, rendering the small configuration appropriate for balancing responsiveness with detection fidelity. Desktop GPUs and server-class accelerators provide sufficient resources to support medium through extra-large configurations, with the specific choice determined by latency requirements and operational modes. For batch processing of archived inspection imagery where throughput rather than latency governs efficiency, the extra-large configuration maximizes detection accuracy without deployment constraints.
@@ 3.7 Modality Ablation Study
Table 5 presents the modality ablation results, quantifying the individual and combined contributions of RGB and depth inputs to detection performance.
The RGB-only configuration substantially outperforms the depth-only baseline across all metrics, with mAP50 of 0.193 versus 0.151 and APs of 0.073 versus 0.046. This disparity confirms that appearance information remains the primary discriminative source for transmission-line defect detection, as RGB imagery captures texture variations, color anomalies, and fine-grained surface patterns that depth sensors cannot reliably resolve. The depth-only setting nonetheless achieves non-trivial performance, indicating that geometric cues alone provide meaningful supervision for defect localization despite inherent sensor limitations such as quantization noise and edge artifacts.
Integrating both modalities through CMAFNet yields consistent and substantial improvements over either single-modality baseline. The fused configuration achieves mAP50 of 0.244, representing gains of 26.4% over RGB-only and 61.6% over depth-only settings. More notably, APs increases from 0.073 to 0.089, demonstrating a 21.9% relative improvement in small-object detection—the predominant challenge in transmission-line inspection where 94.5% of defect instances qualify as small objects. The recall metrics exhibit similar trends: insulator contamination recall rises from 0.367 (RGB) to 0.477 (RGB+Depth), and conductor strand breakage recall improves from 0.310 to 0.365. These category-specific gains indicate that depth information provides complementary boundary and surface discontinuity cues that enhance detection of geometrically subtle defects exhibiting low chromatic contrast against cluttered backgrounds.
The precision-recall trade-off merits attention. While RGB-only achieves the highest precision (0.677), it exhibits low recall (0.187), indicating a conservative detection strategy that misses numerous true positives. The fused model accepts a moderate precision reduction (0.452) in exchange for substantially higher recall (0.232), yielding a net improvement in overall detection performance as reflected in mAP metrics. This trade-off aligns with practical inspection requirements where false negatives carry greater risk than false positives.
From a computational perspective, the multimodal configuration introduces marginal overhead—5.5M parameters and 12.9 GFLOPs compared to 4.9M and 12.4 GFLOPs for single-modality baselines—while maintaining real-time inference at 222.3 FPS. These results collectively validate both the effectiveness of depth as a complementary modality and the efficiency of the proposed cross-modal alignment and fusion strategy.
Table 5: Modality ablation of CMAFNet on the TL-RGBD dataset. Best results for each metric are highlighted in bold.